OpenStudy (anonymous):

Find the slope of the line that is parallel to the line that goes through the points (-7,4) and (5,-4)

OpenStudy (luigi0210):

First find the slope. Parallel meaning the same so you don't need to do anything to it. Then plug that into the point-slope formula \(\large y-y_{1}=m(x-x_{1})\)

OpenStudy (anonymous):

use (y2-y1) over (x2-x1)

OpenStudy (luigi0210):

To find the slope, yes.

OpenStudy (anonymous):

okay once i did that i plug it into point-slpoe formula -4-4=-8/12(5-(-7))

OpenStudy (luigi0210):

-8/12? Yea, that'll be your slope, m. Now just plug that back into the point-slope formula.

OpenStudy (anonymous):

i got -8 =-8

OpenStudy (luigi0210):

o.O How? It's suppose to be an equation.

OpenStudy (anonymous):

im not sure. could you show me

OpenStudy (luigi0210):

We will use (-7, 4). Plug that in with the slope: \(\large y-4=\frac{-2}{3}(x+7)\) Now just put that into slope intercept form.

OpenStudy (anonymous):

im not sure how to do that

OpenStudy (luigi0210):

Simplify it.

OpenStudy (anonymous):

2/3